OpenStudy (anonymous):

hookes law. the distance when a spring is stretched by a hanging object varies directly as the weight w of the object if the distance is 70cm when the weight is 3kg what is the distance when the weight is 9kg

OpenStudy (anonymous):

d=kw k=d/w =70/3 d=70/3*9=210

OpenStudy (anonymous):

70/3kg=x/9kg 70*9=3x 630=3x x=210. DANGIT

OpenStudy (anonymous):

your fast uzma LOL